Alex is 6 yrs. older than Frank. the sum of their ages is 50. find alex and frank age. show solution.
Found 2 solutions by ewatrrr, Earlsdon:
Answer by ewatrrr(24785)


Hi
Alex is 6 yrs. older than Frank
Let x and (x+6) represent Frank's and Alex's ages respectively
question states***
x +(x+6) = 50
solving for x
2x + 6 = 50
2x = 44
x = 22, Franks age. Alex is 28. (22 + 6)
CHECKING our Answer
22 + 28 = 50

Answer by Earlsdon(6294)

Let A = Alex's age and F = Frank's age.
A = F+6 "Alex is 6 yrs. older than Frank."
A+F = 50 ""The sum of their ages is 50." Substitute for A, A = F+6 and solve for F.
(F+6)+F = 50
2F+6 = 50
2F = 44
F = 22 and A = 22+6 = 28.
Alex is 28 years old and Frank is 22 years old.